### v1

@@ II.2 Nano-imaging of moiré exciton using photoemission microscopes
Figure 4-(b) shows the momentum-integrated energy distribution of the tr-ARPES signal in WSe-MoS heterostructures Bange et al. (2024). The highest contribution at early delay times occurs at around 1.7 eV (orange dashed line) corresponding to the energy of the optically excited KK intralayer excitons in the WSe layer. Interestingly, one observes a signal transition to a second long-lived peak at the lower energy of about 1.1 eV (red dashed line) occurring on a femtosecond timescale. This peak has been identified as the photoemitted electronic contribution of the KK interlayer exciton. Here, the first/second capital letter denotes the position of the hole/electron in the reciprocal space (K, valley), while the subindex stands for the layer (W or Mo). To support this assignment, the same measurement was performed on a WSe monolayer. No spectral weight is observed in the momentum and energy space region associated with the interlayer excitons.
@@
The observed charge transfer dynamics, Fig. 4(c), involves a phonon-driven two-step scattering process Meneghini et al. (2022). The optically excited intralayer KK excitons in the WSe layer cannot efficiently scatter directly to the energetically lowest KK interlayer exciton states (hole at the K valley of the WSe layer and electron in the K valley of the MoSe) [Fig. 3-(b)]. The KK excitons rather scatter first with phonons to the momentum-dark K excitons, where the electrons at the valley are delocalized over both layers. In a second step, these hybridized excitons scatter into the KK interlayer excitons, where the electron has been transferred to the MoSe layer. The direct scattering process is prevented by the weak wave function overlaps at the K valley. In contrast, the electronic wave function at the valley has large contributions at the outer selenium atoms, resulting in much more efficient hybridization of K states Merkl et al. (2020).
@@
Tr-ARPES has also revealed the hole transfer process mechanism across the WSe-MoS interface Meneghini et al. (2023). The K-K intralayer excitons in the MoS layer can be resonantly excited at 1.9 eV (Fig. 4-(d)). However, this still leads to a partial electron transfer due to the off-resonant excitation of the energetically lower K-K excitons (1.7 eV) in the WSe layer. To distinguish the impact of the electron and hole transfer, the temporally resolved formation of K-K interlayer excitons have been studied for both scenarios, i.e. excitation at 1.7 eV (only electron transfer expected) and excitation at 1.9 eV (electron and hole transfer expected). For the electron-only transfer process, the interlayer excitons are formed quickly ( 35 fs), and the ARPES signal saturates on the sub-200 fs timescale (Fig. 4-(d)). In contrast, the joint electron- and hole-transfer is found to be much slower, with a formation time of approximately 3 ps. This can be traced back to the different relaxation pathways of optically excited K-K and K-K excitons. The degree of the hybridization of the intermediate hybrid states and the energetic offset to the initial and final states determine the efficiency of the charge transfer process. While the first scatter via hybridized -K excitons, the latter scatter via hybridized K- states towards to energetically lowest K-K interlayer excitons.
Assuming that ARPES only provides information on the ejected electrons, one would not expect any difference whether the measured photoelectrons stem from the break-up of the intralayer K-K or interlayer K-K excitons as the electron remains in the MoSe layer, while the hole transfers across the interface (Fig. 4-(d)). Interestingly, one finds a non-intuitive blue shift of about 170 meV of the ARPES signal for increasing delay times (Fig. 3-(e)). During this time, the optically excited intralayer excitons scatter into the energetically lower interlayer states, so one would rather naively expect a red-shift in the ARPES signal. Auger processes and energy renormalization could be excluded as possible explanations for the unexpected blue-shift. The unintuitive shift towards higher energies is a hallmark of the correlated nature of the electron-hole pairs.
@@
Typically, tr-ARPES experiments do not spatially resolve the carrier dynamics, but they rather average over large sample areas of typically 10m or more. Thus, they cannot access spatial inhomogeneities, dielectric disorder, strain gradients, or local lattice reconstructions. Recently, ultrafast dark field momentum microscopy has been introduced as a new technique for simultaneous nano-imaging and nano-spectroscopy of materialsSchmitt et al. (2023). The technique has a temporal resolution of 55 fs and a spatial resolution of 500 nm and has been applied to study the spatial heterogeneity in a WSe-MoS heterostructure as well as its impact on the energy and the formation time of interlayer excitons. Interestingly, distinct spatial heterogeneity is found even in seemingly flat areas of the sample exhibiting high-quality photoemission spectra. Positioning the dark-field aperture on the momenta of the interlayer excitons provides access to the spatially and spectrally resolved evolution of the interlayer exciton energy. The nanoscale variation of the latter as well as its difference to the energy of the optically excited intralayer exciton is shown in spatial heat maps in Fig. 4-(f). Interestingly, the experiment reveals a correlation between the interlayer exciton formation time and the energy difference , i.e. gets larger with the increasing (Fig.4-(g)). This observation is qualitatively confirmed by the many-particle theory (inset), which could trace this correlation back to nanoscale variations of the strength of interlayer hybridization in the WSe-MoS bilayers. This results in a changed efficiency for the phonon-driven charge transfer to the intermediate hybridized states and thus has a direct impact on the formation time of interlayer excitons.
@@ II.3 Atomic relaxation and its influence on moiré excitons
The emergence of two-dimensional van der Waals materials has raised the intriguing idea of mixing and matching layers of different materials to realize novel properties in a given stack very early on. With the rush to build new Lego-like heterostructures in which to combine the vastly different electronic and magnetic ground states Geim and Grigorieva (2013), two fundamental real space properties of van der Waals materials have been predominantly ignored: (i) their lattice constants are extremely dependent on local changes of composition and (ii) the atomistic positions in van der Waals materials are not rigid, but rather follow a membrane-like behavior. The second point becomes most prominent in large moiré unit cells formed by small twist angles between adjacent layers or by small lattice constant mismatches in heterostructures. In this scenario, the relaxation of atomistic positions in and out of the plane can minimize the energy of the system, and the reconstructed atomistic registry might be very different from that expected for unrelaxed, rigid heterobilayers with severe consequences for electronic, excitonic and mechanical properties Carr et al. (2018a); Halbertal et al. (2022); Susarla et al. (2022). Indeed, today, it is known that atomic corrugation plays a crucial role in the celebrated flat bands and, in particular, the energy gaps of twisted bilayer graphene at the magic angle Uchida et al. (2014); Lucignano et al. (2019). Recently, even the role of moiré phonons, as vibrational stacking re-arrangements, have been highlighted as an unconventional way to boost exciton diffusion Rossi et al. (2023). Thus, it is fundamental to characterize experimentally the local stacking arrangements. Piezoresponse force microscopy (PPM) is an extension of atomic force microscopy that has emerged as a universal and easily implemented technique to map the real space variation of the moiré potential McGilly et al. (2020). PPM measures the amplitude and phase of the electromechanical response of a material to local deformations induced by an AC bias between its surface and the AFM tip. PPM maps the moiré superlattice by recording the changes in the flexoelectric response of heterobilayers associated with the inherent strain gradients of moiré heterostructures. This manifestation of the moiré potential as polar domain walls in a non-polar background is irrespective of the inversion symmetry of the constituent mono-layers McGilly et al. (2020).
@@
While scanning tunneling microscopy and spectroscopy experiments (STM/S) lack the momentum resolution to resolve the reciprocal space information of the electronic band structure of moiré heterobilayers, STM/S has emerged as the only techniques with the required atomic and energy resolution to observe changes to the excitonic landscape due to atomic relaxation. For the STM/STS technique and its application to quantum materials, we refer the reader to other previous reviews Bian et al. (2021); Yin et al. (2021). Early STM results on moiré heterostructures including Ref. Zhang et al. (2017, 2020b); Li et al. (2021a) were reviewed in Ref. Huang et al. (2022). We now discuss examples of the effects of local strain in the electronic structure of moiré heterobilayers. In small-angle heterostructures of WSe/MoSe STM experiments revealed spatially varied moiré period up ranging from to wavelengths over a nm Shabani et al. (2021). The resulting strain associated with the moiréstructure induces systematic large variations on the moiré potential depths and local changes to the near states, see Fig. 5. In twisted bilayers of WS STS revealed the existence of flat bands associated with localized states that cannot be accounted for with first-principle calculations, but that can be recovered when the atomic displacements due to strain are includedMolino et al. (2023). STM has also been deployed to observe changes of more than in the moiré period of heterostructures of MBE grown multilayer PdTe on BiSe (Figure 6) Halbertal et al. (2023).
@@
In studying intra- and interlayer moiré excitons, it is also vital to understand and characterize the precise atomistic registry as it will modify the specific interactions and relaxation pathways. Laser-STM experiments in which STM/STS measurements are performed while a continuous-wave laser shines on the sample, enabling resolving the electronic structure of moiré excitons with sub-nm resolution. Laser-STM experiments in twisted bilayer WS revealed nanoscale modulation in the moiré excitons Li et al. (2024) (Figure 5). The STM topography maps in these heterostructures are characterized by a moiré pattern with a 9 nm period. STS measurements show clear differences in the electronic band structure at B and AB stacking sites reflecting the deep moire potential in twisted WS. Upon turning the laser on, the emerging excitonic photocurrent displays different behavior for both sites. When the bias voltage between tip and sample is set in the region , the photocurrent is concomitantly positive in the B sites and a negative in the AB stacking sites (Figure 5), indicating a dominant electron and hole contribution respectively. This phenomenology was explained by GW-Bethe–Salpeter equation calculations, including the moiré potential and electron-hole Coulomb interactions, which indicate the formation of in-plane charge transfer (ICT) moiré excitons. Thus,the STM tip-position photocurrent can be explained as a result of the lateral separation of electrons and holes in an ICT excitons Li et al. (2024).
@@
Optical pump-probe STM setups with picosecond resolution could enable studying the atomic dependence of exciton dynamics Liang et al. (2023). The challenge in tr-STM measurements is to remove thermal fluctuations associated with the optical pump-induced expansion of the atomic tip and sample, which can result in large artifacts in the tunneling currents. The development of shaken-pulse-pair-excited STM, in which the pump beam is electronically modulated by an external trigger acting as the reference signal for lock-in amplifier detectionMogi et al. (2019); Takeuchi et al. (2002), has enabled the required high-signal-to-noise ratios and time resolution to study non-equilibrium dynamics with atomic resolution Iwaya et al. (2023). Recently, tr-STM has been deployed to study the spatial dependence exciton dynamics in heterobilayers of WS/WSe Mogi et al. (2022) near grain boundaries and strain-induced ripples (see Figure 5). The authors revealed a larger exciton lifetime between the top and bottom of the ripple, suggesting that interaction with the substrate phonon modes can lead to shorter excitonic lifetimes. Similarly, substrate-induced strain near grain boundaries also induced faster excitonic relaxation. The effect diminishes rapidly in space, with measurements nm away from the grain boundary, where long exciton lifetimes are recovered.
@@
Theoretical modeling of the atomic relaxation requires taking the multi-layered nature of vdW heterostructures into account properly, as it has been shown that the strain fields emerging from moiré defects can penetrate deep into a vdW material stack Cook et al. (2022); Halbertal et al. (2023). However, a theoretical description of vdW heterostructures with large moiré periods and/or many layers on the atomic scale is computationally out of reach, and as a consequence, sophisticated coarse-grained models were developed that address atomic relaxation at large length scales. If the moiré lattice spans many unit cells of a single layer, the local atomic stacking registry and in- and out-of-plane strain fields, can be described by generalized forces acting on a stack of membranes (i.e., one models each single layer as one of these membranes) Carr et al. (2020, 2018b); Halbertal et al. (2023); Zhu et al. (2020); Cazeaux et al. (2020); Halbertal et al. (2021). In such membrane models, each layer of the vdW stack is allowed to relax individually. The total energy functional is then a sum of the generalized stacking fault energy, which accounts for lattice mismatch at any of the stack’s interfaces, and the elastic energy, which accounts for the in- and out-of-plane strain of each of the membranes. The parameters for both the elastic and the stacking fault energy are readily obtained from ab-initio simulations of the individual constituent layers and interfaces Halbertal et al. (2023). The multi-layered relaxation process in vdW heterostructures can heavily affect electronic and, thus, excitonic properties. Therefore, the theoretical modeling of relaxation, e.g., by the aforementioned membrane models, is an essential ingredient also in describing these properties in moiré structures even far away from the interface Halbertal et al. (2023); Mullan et al. (2023).
@@
These experimental and theoretical studies exemplify that it is essential to understand the relaxation mechanisms of moiré heterostructures toward fully understanding their emergent excitonic properties. In a second step one might consider using long-ranged relaxation even for engineering-by-stacking purposes.

### v2

@@ II.2 Real and reciprocal space imaging of moiré excitons using photoemission microscopes
We proceed to review hybrid excitons in moiré heterobilayers. In the following, the first/second capital letter denotes the position of the hole/electron in the reciprocal space (K, valley), while the subindex stands for the layer (W for WS2 and Mo for MoS2). The left panel of Figure 4 (b) shows the momentum-integrated energy distribution of the tr-ARPES signal in WSe2-MoS2 moiré heterostructures Schmitt et al. (2022). The highest contribution at early delay times occurs at around 1.7 eV (orange dashed line), corresponding to the energy of the optically excited KK intralayer excitons in the WSe2 layer. A transition to a second long-lived peak at the lower energy of about 1.1 eV (red dashed line) occurring on a femtosecond timescale is observed. This peak has been identified as the photoemitted electronic contribution of the K-K interlayer hybrid exciton (hole at the K valley of the WSe2 layer and electron in the K valley of the MoS2) [Figure 3(b)]. The same measurement was performed on a WSe2 monolayer to support this assignment, as shown in the right panel of Figure 4 (b). No spectral weight is observed in the momentum and energy space region associated with the interlayer excitons in the monolayer, highlighting the hybrid nature of the exciton in WSe2-MoS2 heterostructures.
@@
The 1D plots in Figure 4 (c) summarize the observed charge transfer dynamics shown in the 2D plots of Figure 4 (b). The optically excited intralayer KK excitons in the WSe2 layer cannot efficiently scatter directly to the energetically lowest KK interlayer exciton states. The KK excitons rather scatter first with phonons to the momentum-dark K- excitons, where the electrons at the valley are delocalized over both layers. In a second step, these hybridized excitons scatter into the KK interlayer excitons, where the electron has been transferred to the MoSe2 layer Meneghini et al. (2022). We note that the direct scattering process is prevented by the weak wave function overlap at the K valley. In contrast, the electronic wave function at the valley has large contributions at the outer selenium atoms, resulting in much more efficient hybridization of K- states Merkl et al. (2020).
@@
Tr-ARPES has also revealed the hole transfer process mechanism across the WSe2-MoS2 interface Meneghini et al. (2023). The K-K intralayer excitons in the MoS2 layer can be resonantly excited at 1.9 eV [Figure 4 (d)]. However, this still leads to a partial electron transfer due to the off-resonant excitation of the energetically lower K-K excitons (1.7 eV) in the WSe2 layer. To distinguish the impact of the electron and hole transfer, the temporally resolved formation of K-K interlayer excitons have been studied for both scenarios, i.e. excitation at 1.7 eV (only electron transfer expected) and excitation at 1.9 eV (electron and hole transfer expected). For the electron-only transfer process, the interlayer excitons are formed quickly ( 35 fs), and the ARPES signal saturates on the sub-200 fs timescale [Figure 4 (d)]. In contrast, the joint electron- and hole-transfer is found to be much slower, with a formation time of approximately 3 ps. This can be traced back to the different relaxation pathways of optically excited K-K and K-K excitons. The degree of the hybridization of the intermediate hybrid states and the energetic offset to the initial and final states determine the efficiency of the charge transfer process. While the first scatter via hybridized -K excitons, the latter scatter via hybridized K- states towards the energetically lowest K-K interlayer excitons. In principle, no difference in the ARPES response is expected whether the measured photoelectrons stem from the break-up of the intralayer K-K or interlayer K-K excitons as in both cases the electron remains in the MoS2 layer while the hole transfers across the interface [Figure 4 (d)]. On the other hand, a blue shift of about 170 meV of the ARPES signal is found at increasing delay times [Figure 4 (e)] instead of the redshift expected in optically excited intralayer excitons due to scattering into the energetically lower interlayer states. Auger processes and energy renormalization could be excluded as possible explanations for the unexpected blue shift. Thus, the shift towards higher energies can only be understood as a hallmark of the correlated nature of the hybridized electron-hole pairs.
@@
Finally, we review ultrafast dark field momentum microscopy, introduced as a new technique for simultaneous nano-imaging and nano-spectroscopy of materials Schmitt et al. (2023). Typically, tr-ARPES experiments do not spatially resolve the carrier dynamics, but they average over large sample areas of typically 10m or more. Thus, they cannot access spatial inhomogeneities, dielectric disorder, strain gradients, or local lattice reconstructions. Ultrafast dark field momentum microscopy has been applied to the study of spatial heterogeneity in a WSe2-MoS2 heterostructure and its impact on the energy and the formation time of hybrid interlayer excitons. Distinct spatial heterogeneity is found even in optically flat areas of the sample exhibiting high-quality photoemission spectra [Figure 4 (f)]. Positioning the dark-field aperture on the momenta of the interlayer excitons provides access to the spatially and spectrally resolved evolution of the interlayer exciton energy. The nanoscale variation of the latter, as well as its difference to the energy of the optically excited intralayer exciton, is shown in spatial heat maps in Figure 4 (f). A correlation was observed between the interlayer exciton formation time and the nanoscale variation of the exciton energy . gets larger with the increasing [Figure 4 (g)]. This observation is qualitatively confirmed by the many-particle theory (inset), which could trace this correlation back to nanoscale variations of the strength of interlayer hybridization in the WSe2-MoS2 bilayers. This study revealed how spatial inhomogeneities directly impact the formation timescale of interlayer excitons by influencing the efficiency of the phonon-driven charge transfer to the intermediate hybridized states.
Despite requiring careful sample preparation, including the need for vacuum annealing to remove surface adsorbates and for electrical isolation from the substrate while avoiding sample charging, tr-ARPES and ultrafast momentum microscopy have emerged as unique tools for understanding the formation of moiré excitons and the band structure of moiré heterostructures.
@@ II.3 Atomic relaxation and its influence on moiré excitons
In recent efforts to build new 2D heterostructures combining distinct electronic and magnetic ground states from different v.d.W. materials Geim and Grigorieva (2013), two fundamental real space properties of these compounds have been predominantly ignored: (i) their lattice constants are extremely dependent on local changes of the composition and (ii) the atomistic positions in v.d.W. materials are not rigid but rather follow a membrane-like behavior. The second point becomes most prominent in large moiré unit cells formed by small twist angles between adjacent layers or by small lattice constant mismatches in heterostructures. In this scenario, the relaxation of atomistic positions in and out of the plane can minimize the energy of the system. Thus, the reconstructed atomistic registry might be very different from that expected for unrelaxed, rigid heterobilayers with severe consequences for electronic, excitonic, and mechanical properties Carr et al. (2018a); Halbertal et al. (2022); Susarla et al. (2022). It has recently been realized that atomic corrugation plays a crucial role in the flat bands and energy gaps of twisted bilayer graphene at the magic angle Uchida et al. (2014); Lucignano et al. (2019). Moreover, vibrational stacking re-arrangements, known as moiré phonons, have been highlighted as an unconventional way to boost exciton diffusion Rossi et al. (2023); Qian et al. (2024). Thus, it is fundamental to characterize experimentally the local stacking arrangements.
Piezoresponse force microscopy (PFM) is an extension of atomic force microscopy (AFM) that has emerged as a universal and easily implemented technique to map the real space variation of the moiré potential McGilly et al. (2020). PFM measures the amplitude and phase of the electromechanical response of a material to local deformations induced by an alternating current (AC) bias between its surface and the AFM tip [Figure 5 (a)]. The resolution figure of merit in a PFM experiment is given by the contact radius of the AFM, which is generally around nm. PFM is a highly versatile tool that can quickly identify the local variations of the moiré potential. When deployed to moiré superlattices, irrespectively of the inversion symmetry of the constituent mono-layers, PFM can track the changes in the flexoelectric response due to inherent strain gradients in the heterostructure leading to polar domain walls in a non-polar background. McGilly et al. (2020). This is exemplified in Figure 5 (c) and (d), where the PFM amplitude and phase change show the domain wall array as a result of the moiré pattern in a twisted bilayer graphene (tBLG) sample [Figure 5 (b)]. Despite the centrosymmetric nature of tBLG, a large PFM response is found at the AA stacking site and domain walls, while AB domains give a negligible response.
Going beyond nm resolution, scanning tunneling microscopy and spectroscopy experiments (STM/S) emerge as the only techniques with the required atomic and energy resolution to observe changes to the excitonic landscape due to atomic relaxation. For an in-depth review of the STM/STS technique and its application to quantum materials, we refer the reader to previous reviews Bian et al. (2021); Yin et al. (2021). Additionally, early STM/S measurement on moiré heterostructures Zhang et al. (2017, 2020b); Li et al. (2021a) were reviewed in Ref. Huang et al. (2022).
@@
We now discuss examples of the effects of local strain in the electronic structure of moiré heterobilayers, as revealed by STM/S. In small-angle heterostructures of WSe2/MoSe2 STM experiments revealed spatially varied moiré period up ranging from to nm wavelengths over a nm2 Shabani et al. (2021) [Figure 6 (a)]. The resulting strain associated with the moiré structure induces systematic large variations on the moiré potential depths that can be recovered by calculations considering the effects of generalized forces acting on a stack of membranes Carr et al. (2018b) [Figure 6 (b)]. Moreover, STS measurements reveal local changes to the near states with the moiré potential wavelength, with the electronic gap narrowing at longer wavelengths [Figure 6 (c)]. Additionally, in twisted bilayers of WS2 STS revealed the existence of flat bands associated with localized states that cannot be accounted for with first-principle calculations but that can be recovered when the atomic displacements due to strain are included Molino et al. (2023). STM has also been deployed to observe changes of more than in the moiré period of heterostructures of molecular-beam epitaxy (MBE) grown multilayer PdTe2 on Bi2Se3 [Figure 7 (b)] Halbertal et al. (2023).
@@
In studying intra- and interlayer moiré excitons, it is also vital to understand and characterize the precise atomistic registry as it will modify the specific interactions and relaxation pathways. Laser-STM experiments in which STM/STS measurements are performed while a continuous-wave laser shines on the sample, enabling resolving the electronic structure of moiré excitons with sub-nm resolution. Laser-STM experiments in twisted bilayer WS2 revealed nanoscale modulation in the moiré excitons Li et al. (2024). The STM topography maps in these heterostructures are characterized by a moiré pattern with a 9 nm period [Figure 6 (e)]. STS measurements show clear differences in the electronic band structure at BW/W, direct stacking of W atoms of different layers, and AB stacking sites reflecting the deep moiré potential in twisted WS2 [Figure 6 (d)]. Upon turning the laser on, the emerging excitonic photocurrent displays different behavior for both sites. When the bias voltage between tip and sample is set in the region , the photocurrent is concomitantly positive in the BW/W sites and a negative in the AB stacking sites [Figure 6 (f)], indicating a dominant electron and hole contribution respectively. This phenomenology was explained by GW-Bethe–Salpeter equation calculations, including the moiré potential and electron-hole Coulomb interactions, which indicate the formation of in-plane charge transfer (ICT) moiré excitons. Thus, the STM tip-position photocurrent can be explained as a result of the lateral separation of electrons and holes in an ICT exciton [Figure 6 (g)] Li et al. (2024).
@@
Optical pump-probe STM setups with picosecond resolution could enable studying the atomic dependence of exciton dynamics Liang et al. (2023). The challenge in tr-STM measurements is to remove thermal fluctuations associated with the optical pump-induced expansion of the atomic tip and sample, which can result in large artifacts in the tunneling currents. The development of shaken-pulse-pair-excited STM, in which the pump beam is electronically modulated by an external trigger acting as the reference signal for lock-in amplifier detection Mogi et al. (2019); Takeuchi et al. (2002), has enabled the required high-signal-to-noise ratios and time resolution to study non-equilibrium dynamics with atomic resolution Iwaya et al. (2023). Recently, tr-STM has been deployed to study the spatial dependence exciton dynamics in heterobilayers of WS2/WSe2 Mogi et al. (2022) near grain boundaries and strain-induced ripples [Figure 6 (h)]. The authors revealed an almost fivefold enhancement of the exciton lifetime between the top and bottom of the ripple, suggesting that interaction with the substrate phonon modes can lead to shorter excitonic lifetimes [Figure 6 (i) - (j)]. Similarly, substrate-induced strain near grain boundaries also leads to three times faster excitonic relaxation than at the grain boundary Mogi et al. (2022). However, the effect diminishes rapidly in space, nm away from the grain boundary, where long exciton lifetimes are recovered Mogi et al. (2022).
@@
Theoretical modeling of the atomic relaxation requires properly considering the multi-layered nature of v.d.W. heterostructures. It has been shown that the strain fields emerging from moiré defects can penetrate deep into a v.d.W. material stack Cook et al. (2022); Halbertal et al. (2023). This departs from the Lego-like interpretation of v.d.W. heterostructures towards a continuous layer-dependent interpretation of defects and strain [Figure 7 (a)]. However, a theoretical description of v.d.W. heterostructures with large moiré periods and/or many layers on the atomic scale is computationally out of reach. As a consequence, sophisticated coarse-grained models were developed that address atomic relaxation at large length scales. When the moiré lattice spans many unit cells of a single layer, the local atomic stacking registry and in- and out-of-plane strain fields can be modeled by considering each single layer as a membrane Carr et al. (2020, 2018b); Halbertal et al. (2023); Zhu et al. (2020); Cazeaux et al. (2020); Halbertal et al. (2021). In such membrane models, each layer of the v.d.W. stack is allowed to relax individually. The total energy functional is calculated as a sum of the generalized stacking fault energy, which accounts for lattice mismatch at stack interfaces, and the elastic energy of each membrane, which accounts for the in- and out-of-plane strain fields. The parameters for both the elastic and the stacking fault energy are readily obtained from ab-initio simulations of the individual constituent layers and interfaces Halbertal et al. (2023). The multi-layered relaxation process in v.d.W. heterostructures can heavily affect electronic and, thus, excitonic properties. Therefore, the theoretical modeling of relaxation, e.g., by the aforementioned membrane models, is an essential ingredient in describing these properties in moiré structures even far away from the interface Halbertal et al. (2023) or in novel interfaces between two three dimensional crystals Mullan et al. (2023) .
@@
These experimental and theoretical studies exemplify that it is essential to understand the atomic relaxation mechanisms of moiré heterostructures to fully understand their emergent excitonic properties. In a second step, one might consider using long-ranged relaxation even for engineering-by-stacking purposes.